OpenStudy (anonymous):

a^8/3 x a^5/6=11/6 IS THIS RIGHT? Simplify your answer.Type exponential notation with rational exponents

OpenStudy (unklerhaukus):

\[x^n\times x^m=x^{n+m}\] \[a^{8/3} \times a^{5/6}=a^{8/3+5/6}\\\qquad\qquad=a^{16/6+5/6}\\\qquad\qquad=a^{(16+5)/6}\]